OpenStudy (anonymous):

suppose a triangle has 3 4 and 6. Which of the following must be true?

OpenStudy (anonymous):

What are the options?

Parth (parthkohli):

Well at least they can be the sides of a triangle :)

OpenStudy (unklerhaukus):

all the angles are different

Parth (parthkohli):

If you wanna check if it is an acute, obtuse, or right triangle check these: ACUTE: \( \color{Black}{\Rightarrow a^2 + b^2 > c^2}\) OBTUSE: \( \color{Black}{\Rightarrow a^2 + b^2 < c^2 }\) RIGHT: \( \color{Black}{\Rightarrow a^2 + b^2 = c^2 }\) c is the longest side.

OpenStudy (anonymous):

well the options are 1. The triangle in question is not a right triangle 2.The triangle in question is a right triangle 3.the triangle in question May or may not be a right triangle

Parth (parthkohli):

Awwyea I was correct

Parth (parthkohli):

Check if \(3^2 + 4^2 = 6^2\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3^2 + 4^2 =/</> 6^2 9 + 16 =/</> 36 25 =/</> 36 25 < 36 25 is less than 36, so the triangle is obtuse. Option 1 is correct.

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

another question

OpenStudy (anonymous):

suppose a triangle has two sides of length 2 and 5 and the angle between these two sides is pi/3. what is the length of the third side of the triangle